OpenStudy (amberparrishxo):

WILL FAN AND MEDAL! :D Solve for x. Express the answer in simplest form.

OpenStudy (legends):

whats te quesion:)

OpenStudy (amberparrishxo):



OpenStudy (amberparrishxo):

i got 5.5 = x but it has to be in a fraction form.

OpenStudy (anonymous):

\(\dfrac{1}{7} = \dfrac{3}{4}x\) Is it this? \(\dfrac{1}{7} = \dfrac{3}{4x} - 1\) Or this?

OpenStudy (anonymous):

Sorry I meant\(\dfrac{1}{7} = \dfrac{3}{4}x - 1\) For the first one

OpenStudy (amberparrishxo):

Niether. the equation is how it is draw above, with the +1 next to the 4x under 3.

OpenStudy (anonymous):

Oh, so like this? \(\dfrac{1}{7} = \dfrac{3}{4x - 1}\)

OpenStudy (amberparrishxo):

Yes :)

OpenStudy (anonymous):

Oh, okay, well 5.5 is correct.. 0.5 is the same thing as 5/10 so we have: \(5 \dfrac{5}{10}\) Or if you want it as an improper fraction, we multiply the denominator(10) to the whole number(5) and add it to the numerator(5) and we keep the denominator the same. So we have \(5 \dfrac{5}{10} \rightarrow \dfrac{10 \times 5 + 5}{10} \rightarrow \dfrac{55}{10} \rightarrow \dfrac{11}{2}\)

OpenStudy (anonymous):

So your answer is either 5 5/10 in mixed number form, or 11/2 in improper fraction form.

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\sf\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!!!}}}\) You can give medals by clicking 'Best Response'.